OpenStudy (anonymous):

the denominator of a fraction is three more than twice the numerator. If both the numerator and denominator are decreased by three, the simplified result is 5 ___ 13. Find the original fraction

OpenStudy (radar):

Let x = the numerator 2x+3 = the denominator according to the problem : x-3 5 ------- = ---- 2x+3-3 13 cross multiply 13x-39= 10x 3x=39 x=13 numerator 2x+3=26+3=29 the denominator Original fraction is: 13/29

OpenStudy (radar):

Please review and ask any question for clarification. I gotta run but will return when I am notified.